# Turbulence to Order Transitions in Activity Patterned Active Nematics

### v1

@@ Abstract
We numerically study two-dimensional active nematics with periodic activity patterning. For stripes of activity, we observe a transition from two-dimensional to one-dimensional active turbulence as the maximum active force and distance between activity stripes increases, followed by a transition to stable vortices ordered antiferromagnetically along the stripes and ferromagnetically transverse to the stripes. If the stripes are replaced by columns of activity circles, the vortex ordering is lost. A triangular lattice of activity circles can induce the formation of two-dimensional active turbulence. Our results provide a mechanism for inducing non-equilibrium phase transitions in active nematics using activity inhomogeneity, which can be exploited to create activity patterned ordered phases.
@@
In non-equilibrium active fluids, energy consumed at the microscale generates macroscopic flows [1, 2]. Active nematics, where the constituents and forces are also endowed with apolar anisotropy, appear in cytoskeletal filament suspensions with molecular motors [3, 4, 5, 6], motile bacteria dispersed in passive nematic liquid crystals [7, 8], cellular tissues [9, 10], colonies of elongated soil bacteria [11], and acoustically actuated lyotropic chromonic liquid crystals [12]. In large, unconfined active nematics, the system exhibits “active turbulence,” where the flows are chaotic and topological defect pairs spontaneously unbind [13, 14, 15, 16, 17, 18]. There has been growing interest in developing methods to control the flows and phase behavior of active nematics by changing the substrate friction or external boundaries [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. A particularly promising approach is activity patterning, recently proven achievable using light activated motor proteins dispersed in cytoskeletal filament suspensions [37, 38, 39, 40]. Recent theoretical and numerical studies indicate that activity gradients can separate and orient topological defects and induce multiple active regimes [41, 29, 31, 34].
@@
Here we show that introducing periodic active and inactive stripes or spots can produce transitions from two-dimensional (2D) to one-dimensional (1D) active turbulence or a stable vortical flow state as the activity strength and distance between stripes are varied. Surprisingly, we find that the transitions to turbulence shift to higher active region densities for higher activity. As the active regions become less dense, an ordered vortex state is more likely to appear, with antiferromagnetic ordering along the active channels and ferromagnetic ordering transverse to them. Although vortices still appear if the activity stripes are replaced by columns of activity spots, the ordering of the vortices is lost. Our results provide a new mechanism for stabilizing ordered vortices in active nematics and controlling transitions to and from different active turbulence regimes. Crucially, we find that the activity does not need to be homogeneously distributed in order to produce active turbulence, and that the density of activity itself can control the phase behavior of an active nematic system.
Methods– We model an active nematic using the Beris-Edwards equations for the tensor order parameter $\mathbf{Q}$ [42, 43] and a modified Stokes equation for the incompressible fluid velocity $\mathbf{v}$ [44, 6]:
@@
$\frac{\partial\mathbf{Q}}{\partial t}+\left(\mathbf{v}\cdot\nabla\right)\mathbf{Q}$ $=\mathbf{S}-\frac{\delta F}{\delta\mathbf{Q}}$ (1)
@@
$\ell_{h}^{2}\nabla^{2}\mathbf{v}-\mathbf{v}$ $=\nabla p+(1+\ell_{h}^{2})\nabla\cdot\left(\alpha\mathbf{Q}\right)$ (2)
$\nabla\cdot\mathbf{v}$ $=0.$ (3)
Here $\mathbf{S}$ is a generalized tensor advection, $F$ is a Landau-de Gennes free energy, $\ell_{h}$ is the hydrodynamic screening length characterized by the ratio of fluid viscosity and substrate friction, $p$ is the fluid pressure which is determined from the incompressibility constraint, and $\alpha=\alpha(x,y)$ is the inhomogeneous strength of active forces. All lengths are measured in units of the nematic correlation length and times in units of the nematic relaxation time [45].
To solve Eqs. (1) and (2), we use the finite element method from the C++/Matlab package FELICITY [46]. We consider rectangular domains of various sizes with different patterns of activity $\alpha(x,y)$ set to a maximum (minimum) value of $\alpha=\alpha_{0}$ ($\alpha=0$), and use periodic boundary conditions in all cases. Figure 1(a) shows a system of activity stripes oriented along the $y$ direction, where the active (inactive) regions are of length $\ell_{s}$ ($\ell_{i}$). We fix $\ell_{s}=\ell_{h}=10$ so that the screening length is commensurate with the active stripe length, causing the active stripes to behave as channels of active nematics that interact across an inactive region. We also consider active circles of radius $R=10/\sqrt{\pi}$ arranged in either a triangular lattice [Fig. 1(b)] with varied lattice constant $a$, or vertical lines [Fig. 1(c)] with varied inactive length $\ell_{i}$.
@@
Figure 1: Height maps of activity profiles showing activity level $\alpha$. (a) Active stripes with active region width $\ell_{s}$ and inactive region width $\ell_{i}$. (b) Active circles of radius $R$ arranged in a triangular lattice with lattice constant $a$. (c) Active circles of radius $R$ arranged along vertical lines with inactive width $\ell_{i}$.
Results– We find two transitions in the active stripe array as we vary $\ell_{i}$ and $\alpha_{0}$. For small $\ell_{i}$ and $\alpha_{0}$, the behavior qualitatively resembles the active turbulence that would appear if the activity were homogeneous [4, 47, 6, 48], as shown by the plot of the time-averaged normalized vorticity $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ and velocity $\langle\mathbf{v}\rangle$ fields for a system with $\ell_{i}=6$ and $\alpha_{0}=0.7$ in Fig. 2(a). Both quantities have a value close to 00 throughout the domain, and there is no correlation between the active regions and the weak vortical patterns or the time-averaged topological defect density [49, 50, 45]. When either $\alpha_{0}$ or $\ell_{i}$ is increased, as shown in Fig. 2(b) at $\ell_{i}=16$ and $\alpha_{0}=0.7$ and in Fig. 2(b) at $\ell_{i}=6$ and $\alpha_{0}=1.2$, a charge separating behavior arises. Here $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ becomes strongly nonzero only in the active regions while $|\langle\mathbf{v}\rangle|$ is suppressed in the active regions, so that steady flows occur only in the inactive regions between the active stripes. The time-averaged topological defect density separates, and positive topological defects occupy the active regions while negative topological defects remain in the inactive regions [45]. A similar charge separation was recently predicted theoretically and observed numerically for a single stripe of activity [34]. At large $\alpha_{0}$, the system develops stable and sometimes static vortices, as illustrated in Fig. 2(d) at $\ell_{i}=16$ and $\alpha_{0}=1.5$, where $\langle\mathbf{v}\rangle$ has clear vortical signatures. The nematic configuration in this regime consists of two positive defects rotating around one another near the center of a given vortex, while negative defects sit in between the vortices [45].
@@
Figure 2: (a-d) Spatial plots of scaled time-averaged vorticity field $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ (color) with time-averaged velocity field $\langle\mathbf{v}\rangle$ (arrows), where bright (dark) areas indicate $\alpha=\alpha_{0}$ ($\alpha=0$). (a) $\ell_{i}=6$ and $\alpha_{0}=0.7$. (b) $\ell_{i}=16$ and $\alpha_{0}=0.7$. (c) $\ell_{i}=6$ and $\alpha_{0}=1.2$. (d) $\ell_{i}=16$ and $\alpha_{0}=1.5$. Animations of the dynamics of each state are available in the Supplemental Material [45]. (e) Heat map of the logarithm of the excess kurtosis of the vorticity distribution, $\log|K|$, over the whole domain as a function of $\alpha_{0}$ vs $\ell_{i}$. (f) Heat map of $\log|K|$ only measured over active regions as a function of $\alpha_{0}$ vs $\ell_{i}$. Dashed line: points from panel (e) where $\log|K|=0$. Solid line: points from panel (f) where $\log|K|=0$. Letters: locations at which images in panels (a-d) were obtained.
To quantify the transitions, we compute the excess kurtosis of the vorticity distributions $P(\omega)$:
$K[\mathcal{P}(\omega)]\equiv\frac{\mu_{4}}{\sigma^{4}}-3$ (4)
@@
where $\mu_{4}$ is the fourth moment of $P(\omega)$ and $\sigma$ is the standard deviation of $P(\omega)$. As a statistical tool, $K$ characterizes the size of tails of a distribution, with $K=0$ for a Gaussian distribution and $K<0$ ($K>0$) for distributions with smaller (larger) tails. We use $K$ as a measure for active turbulence since it has been shown both computationally and experimentally that active turbulence in active nematics is associated with Gaussian vorticity distributions [51, 48, 31]. In Fig. 2(e) we plot a heat map of $\log|K|$ measured over the entire spatial domain as a function of $\alpha_{0}$ versus $\ell_{i}$. The dashed line indicates points where $\log|K|=0$, and we define the $\log|K|<0$ region below this line to be in 2D active turbulence. Next, we perform a measurement of $K[\mathcal{P}(\omega|\alpha>0.5\alpha_{0})]$, in which we restrict the domain to only the active portions of the sample and exclude the inactive areas. We plot the resulting heat map of $\log|K|$ as a function of $\alpha_{0}$ versus $\ell_{i}$ in Fig. 2(f). The solid line indicates points for which $\log|K|=0$. In the window between the solid and dashed lines, where the active regions produce a turbulent signature but the inactive regions do not, 1D active turbulence has emerged along each active stripe, but turbulent behavior is suppressed in the regions between the stripes. Thus, the two transitions in the system are between 2D and 1D active turbulence, and between 1D turbulence and a vortex state.

### v2

@@ Abstract
We numerically study two-dimensional active nematics with periodic activity patterning. For stripes of activity, we observe a transition from two-dimensional to one-dimensional active turbulence as the maximum active force and distance between activity stripes increases, followed by a transition to stable vortices ordered antiferromagnetically along the stripes and ferromagnetically transverse to the stripes. By comparing to a triangular lattice of activity circles, we find that transitions to two-dimensional active turbulence emerge from interplays between the active length scale and activity density, independent of the patterning geometry. The vortex ordering, on the other hand, is highly sensitive to patterning geometry, which we show by comparing the activity stripes to columns of activity circles, where the vortex ordering is lost. Our results provide a mechanism for inducing non-equilibrium phase transitions in active nematics using activity inhomogeneity, which can be further exploited to create activity patterned ordered phases.
I Introduction
@@
In non-equilibrium active fluids, energy consumed at the microscale generates macroscopic flows [1, 2]. Active nematics, where the constituents and forces are also endowed with apolar anisotropy, appear in cytoskeletal filament suspensions with molecular motors [3, 4, 5, 6], motile bacteria dispersed in passive nematic liquid crystals [7, 8], cellular tissues [9, 10], colonies of elongated soil bacteria [11], and acoustically actuated lyotropic chromonic liquid crystals [12]. In large, unconfined active nematics, the system exhibits “active turbulence,” where the flows are chaotic and topological defect pairs spontaneously unbind [13, 14, 15, 16, 17, 18]. There has been growing interest in developing methods to control the flows and phase behavior of active nematics by changing the substrate friction or external boundaries [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. A particularly promising approach is activity patterning, recently proven achievable using light activated motor proteins dispersed in cytoskeletal filament suspensions [37, 38, 39, 40]. Recent theoretical and numerical studies indicate that activity gradients can separate and orient topological defects and induce multiple active regimes [41, 29, 31, 34, 42, 43].
Active-passive interfaces and emulsions of passive droplets immersed in active nematics have been recently studied, revealing interesting couplings between the passive and active phases that induce novel flow states in the system [44, 45, 46, 47, 48, 49]. In these studies, however, the passive system is typically isotropic, so there is no coupling between the active and inactive regions apart from the interface. A recent study of spatially varying activity patterning focused only on separated, inactive, isotropic regions with a connected active nematic region [31]. In that study, the inactive regions acted similarly to obstacles in the active nematic [27, 32, 33, 35, 36], rather than inducing a coupling between the active and passive regions of the nematic. Here we focus on two-dimensional nematics with active regions separated by inactive, but still nematic regions, which allow for coupling between the active and elastic forces over larger length scales. This is particularly amenable to experimental implementation using the light-activated motor protein system [37, 38, 39, 40], but also has fundamental importance in understanding the effect of inhomogeneous activity on active systems in general. In real active fluids, the energy consumed at the microscale, be it chemical or mechanical, is not homogeneously distributed, so understanding the effects of inhomogeneous activity is important for interpreting the results of experiments.
@@
Here we study inhomogeneous activity patterning by introducing periodic active and inactive stripes or circles. In Sec. II we delineate our model and numerical methods, and describe the quantitative measures that we use to analyze our results. In Sec. III we present results for a system of periodic activity stripes and show that such patterning can produce transitions from two-dimensional (2D) to one-dimensional (1D) active turbulence or a stable vortical flow state as the activity strength and distance between stripes are varied. Surprisingly, we find that the transitions to turbulence shift to higher active region densities for higher activity. We confirm this by comparing to a lattice of active circles, and show that the interplay between the active length scale and characteristic activity patterning length result in the turbulence transition. In Sec. IV we examine the vortical flow state in the activity stripe system and show that the vortices order antiferromagnetically along the active stripes and ferromagnetically transverse to them, with stronger ordering as the stripes get further apart. We also show that although vortices still appear if the activity stripes are replaced by columns of activity circles, the ordering of the vortices is lost. Finally, in Sec. V we summarize our results and suggest directions for future studies. Our results provide a new mechanism for stabilizing ordered vortices in active nematics and controlling transitions to and from different active turbulence regimes. Crucially, we find that the activity does not need to be homogeneously distributed in order to produce active turbulence, and that the density of activity itself can control the phase behavior of an active nematic system.
II Methods
II.1 Active nematic model with spatially varying activity
We model an active nematic using the Beris-Edwards equations for the two-dimensional tensor order parameter $\mathbf{Q}=S\left[\mathbf{n}\otimes\mathbf{n}-(1/2)\mathbf{I}\right]$, where $S$ is the degree of local order, $\mathbf{n}$ is the local orientation of the nematic (the director), and $\mathbf{I}$ is the $2\times 2$ identity matrix, in the presence of a fluid flow $\mathbf{v}$ [50, 51]:
@@ II.1 Active nematic model with spatially varying activity
$\frac{\partial\mathbf{Q}}{\partial t}+\left(\mathbf{v}\cdot\nabla\right)\mathbf{Q}=\mathbf{S}-\frac{1}{\gamma}\left[\frac{\delta F}{\delta\mathbf{Q}}\right]^{TS}$ (1)
where $\gamma$ is a rotational viscosity, $[\cdot]^{TS}$ denotes the traceless, symmetric part of a tensor,
[…]
is the two-dimensional Landau-de Gennes free energy with material parameters $A$ and $C$ and elastic constant $L$, and
$\mathbf{S}=\left(\lambda\mathbf{E}+\bm{\Omega}\right)\left(\mathbf{Q}+\frac{1}{2}\mathbf{I}\right)+\left(\mathbf{Q}+\frac{1}{2}\mathbf{I}\right)\left(\lambda\mathbf{E}-\bm{\Omega}\right)\\ -2\lambda\left(\mathbf{Q}+\frac{1}{2}\mathbf{I}\right)\left(\nabla\mathbf{v}:\mathbf{Q}\right)$ (3)
is a generalized tensor advection with $\mathbf{E}$ the strain rate tensor, $\bm{\Omega}$ the vorticity tensor, and $\lambda$ a dimensionless parameter that characterizes the tendency of the director to align or tumble under shear flows [51]. The flows in the active nematic are generated by an active stress that arises due to inhomogeneity in the nematic texture. These are modeled with a modified Stokes equation for zero Reynolds number flows [52, 6]:
$\eta\nabla^{2}\mathbf{v}-\Gamma\mathbf{v}=\nabla p+\nabla\cdot\left(\alpha\mathbf{Q}\right)$ (4)
where $\eta$ is the shear viscosity, $\Gamma$ is the coefficient of friction with the substrate, $p$ is the fluid pressure, and $\alpha=\alpha(x,y)$ is the inhomogeneous strength of active forces. We assume that the fluid is incompressible so that $\nabla\cdot\mathbf{v}=0$, and thus the fluid pressure is determined from this constraint. We note that we neglect elastic stresses that may appear in Eq. (4) as they are higher order in $\mathbf{Q}$ or gradients of $\mathbf{Q}$, and it has been shown that the active stress dominates the elastic stresses when they are included in Eq. (4) [13, 53, 54]. We expect the results presented below to hold for models which include elastic stresses, though the quantitative values of critical parameters may differ.
There are three important length scales intrinsic to the model. The nematic correlation length $\xi=\sqrt{\varepsilon L/C}$ is the length scale over which the nematic distorts, where $\varepsilon$ is a dimensionless parameter which controls the size of topological defects. The hydrodynamic theory induces two more length scales: the hydrodynamic screening length $\ell_{h}=\sqrt{\eta/\Gamma}$ is the scale over which vortices decay due to substrate friction, while the active length $\ell_{\alpha}=\sqrt{L/\alpha}$ is the scale over which active forces perturb the underlying nematic order. The active length scale is set by competing elastic and active forces and is related to the average separation of topological defects. In finite sized systems, if the active length is larger than the system size, elastic forces dominate and defect nucleation is suppressed. On the other hand, if the active length is much smaller than the system size, the system will be characterized by active turbulence [55, 21].
We note that, in this study, we modify Eq. (4) by first dividing by $\Gamma$, then multiplying the active force term by $(\eta+\xi^{2}\Gamma)/(\eta+\xi^{2}\Gamma)$. After these manipulations, Eq. (4) may be written as
@@
$\ell_{h}^{2}\nabla^{2}\mathbf{v}-\mathbf{v}=\nabla\left(\frac{p}{\Gamma}\right)+\left(\xi^{2}+\ell_{h}^{2}\right)\nabla\cdot\left(\frac{\alpha}{\eta+\xi^{2}\Gamma}\mathbf{Q}\right).$ (5)
We do this so that the parameter $\alpha/(\eta+\xi^{2}\Gamma)$ is clearly the parameter responsible for the strength of flows [56].
We now introduce a scheme to non-dimensionalize the above equations, and reduce the system to dimensionless parameters. We normalize all lengths by the nematic correlation length $\xi$ and times by the nematic relaxation time $\tau=\gamma/(\xi^{2}C)$. We may then represent dimensionless space and time by
$\mathbf{\tilde{r}}=\frac{\mathbf{r}}{\xi},\,\tilde{t}=\frac{t}{\tau}.$ (6)
Defining
$\mathbf{\tilde{v}}=\frac{\tau}{\xi}\mathbf{v},\,\mathbf{\tilde{S}}=\frac{\mathbf{S}}{\tau},\,\tilde{F}=\frac{F}{\xi^{2}C},\,\tilde{A}=\frac{A}{C}$ (7)
allows us to rewrite Eq. (1) in a dimensionless form:
$\frac{\partial\mathbf{Q}}{\partial\tilde{t}}+\left(\mathbf{\tilde{v}}\cdot\tilde{\nabla}\right)\mathbf{Q}=\mathbf{\tilde{S}}-\left[\frac{\delta\tilde{F}}{\delta\mathbf{Q}}\right]^{TS}$ (8)
where $\tilde{\nabla}$ denotes a derivative with respect to $\mathbf{\tilde{x}}$ and
[…]
Similarly, we define the dimensionless parameters
$\tilde{\ell_{h}}=\frac{\ell_{h}}{\xi},\,\tilde{p}=\frac{\tau p}{\xi^{2}\Gamma},\,\tilde{\alpha}=\frac{\tau\alpha}{\eta+\xi^{2}\Gamma}$ (10)
which allow us to rewrite Eq. (5) as
$\tilde{\ell_{h}}^{2}\tilde{\nabla}^{2}\mathbf{\tilde{v}}-\mathbf{\tilde{v}}=\tilde{\nabla}\tilde{p}+\left(1+\tilde{\ell_{h}}^{2}\right)\tilde{\nabla}\cdot\left(\tilde{\alpha}\mathbf{Q}\right).$ (11)
We note that under this non-dimensionalization scheme we may also define a dimensionless active length $\tilde{\ell_{\alpha}}=\ell_{\alpha}/\xi$ such that $\tilde{\ell_{\alpha}}\propto 1/\sqrt{\varepsilon\tilde{\alpha}}$. For the remainder of the paper we omit the tildes and work in dimensionless units. In particular, all lengths are given in units of the nematic correlation length.
@@
Figure 1: Height maps of activity profiles showing activity level $\alpha$. (a) Active stripes with active region width $\ell_{s}$ and inactive region width $\ell_{i}$. (b) Active circles of radius $R$ arranged in a triangular lattice with lattice constant $a$. (c) Active circles of radius $R$ arranged along vertical columns with inactive width $\ell_{i}$.
To study activity inhomogeneity, we consider rectangular domains of various sizes with different patterns of activity $\alpha(x,y)$ set to a maximum value of $\alpha=\alpha_{0}$ and minimum value of $\alpha=0$. Figure 1(a) shows a system of active stripes oriented along the $y$ direction, where the active (inactive) regions are of length $\ell_{s}$ ($\ell_{i}$) relative to the nematic correlation length. The equation for the stripe pattern is
$\alpha(x,y)=\frac{\alpha_{0}}{2}\sum_{n=1}^{N=6}\left[1+\tanh\left(\frac{\ell_{s}^{2}/4-(x-C_{n})^{2}}{\delta^{2}}\right)\right]$ (12)
where $C_{n}$ give the strip centers which are equally spaced at intervals of $\ell_{s}+\ell_{i}$ and $\delta$ is a parameter that controls the width of the interface between active and inactive regions. We study systems with six activity stripes so that the system size in the $x$ direction is $L_{x}=6(\ell_{i}+\ell_{s})$ while we fix the system size in the $y$ direction to be $L_{y}=8\ell_{h}$.
We also consider active circles of radius $R$ arranged in either a triangular lattice with lattice constant $a$ [Fig. 1(b)] or vertical columns with inactive length spacing $\ell_{i}$ [Fig. 1(c)]. In either case, the equation for the activity is
$\alpha(x,y)=\frac{\alpha_{0}}{2}\sum_{n=1}^{N}\left[1+\tanh\left(\frac{R-|\mathbf{r}-\mathbf{C}_{n}|}{\delta}\right)\right]$ (13)
where $\mathbf{r}=(x,y)$ and $\mathbf{C}_{n}$ is the vector position of the circle centers. For the lattice of activity circles, the number of activity circles varied with the lattice constant $a$ (or area fraction $\nu$) so that the system size in the $x$ direction $L_{x}=\sqrt{N}a$ while the system size in the $y$ direction was $L_{y}=\sqrt{3N}a/2$. For the smallest area fractions we used $N=4$ activity circles, while for the largest area fractions we used $N=100$, choosing the number that kept $L_{x}$ as close to $120$ as possible. For the vertical columns of circles we used a fixed $N=48$ active circles with system sizes $L_{x}=6(\ell_{i}+2R)$ and $L_{y}=16R$.
II.2 Numerical Implementation
To numerically solve Eqs. (8) and (11), we write $\mathbf{Q}$ in a basis for symmetric, traceless matrices, giving two degrees of freedom $q_{1}$ and $q_{2}$. Equation (8) is then rewritten in terms of $q_{1}$ and $q_{2}$. We discretize the equations in space using the finite element Matlab/C++ package FELICITY [57] and write Eqs. (8) and (11) in weak form, using the pressure as a Lagrange multiplier for the incompressibility constraint. We note that spatial derivatives of $\alpha$ also contribute to the flow when $\alpha$ is inhomogeneous. We compute these from Eqs. (12) and (13).
Given a nematic configuration $\mathbf{Q}$, we first compute the corresponding flow velocity from Eq. (11). The resulting velocity is then used to solve for the updated $\mathbf{Q}$ from Eq. (8) using a backwards Euler method with a Newton-Raphson sub-routine to solve the resulting nonlinear equations. For all simulations we use periodic boundary conditions on the rectangular domain, the size of which is determined by the inactive length $\ell_{i}$ or lattice constant $a$. For all simulations we set $A=-2$, $\lambda=1$, $\varepsilon=4$, and the Euler time step $\delta t=0.5$. We also fix the hydrodynamic screening length and the activity patterning length scales to be commensurate, so $\ell_{h}=10$, $\ell_{s}=10$, and $R=10/\sqrt{\pi}$ are not varied in this study. We do this so the size of the active regions will be roughly the size of a single fluid vortex. This is to promote comparison with previous studies of active nematics in channel geometries that produced vortices of roughly channel height [55, 58]. The parameters that we vary are the maximum activity $\alpha_{0}$ and the inactive region lengths $\ell_{i}$ or $a$.
We initialize the system with a random director configuration and simulate in the range of $300$ to $3000$ time steps depending on how long the system takes to reach a stationary configuration or a dynamical steady state. We define a system to be in a dynamical steady state when the time average of measured quantities does not appreciably change if the time averaging is performed over later times.
II.3 Measures
To analyze the behavior of the activity patterned active nematic system, we use several measures involving the spatially varying fields $\mathbf{Q}$ and $\mathbf{v}$. To understand the distribution of topological defects — singular points in the director field around which the director winds by half integer multiples of $2\pi$ — we compute the topological charge density [59, 60]
$D=\frac{1}{2\pi S_{N}^{2}}\varepsilon_{k\ell}\varepsilon_{\mu\nu}\partial_{k}Q_{\mu\alpha}\partial_{\ell}Q_{\nu\alpha}$ (14)
where $S_{N}$ is the value of $S$ in the nematic phase (here $S_{N}=\sqrt{2}$), $\bm{\varepsilon}$ is the two-dimensional fully antisymmetric tensor, $\partial_{k}\equiv\partial/\partial x_{k}$, and summation on repeated indices is assumed. $D$ has the property that it is zero away from defects, positive at the core of $+1/2$ winding defects, and negative at the core of $-1/2$ defects.
We also use the statistics of the vorticity field, $\omega=\left(\nabla\times\mathbf{v}\right)_{z}$, to analyze the state of a system. In particular, we make use of the excess kurtosis of the vorticity distribution, defined by
$K\left[\mathcal{P}(\omega)\right]=\frac{\mu_{4}}{\sigma^{4}}-3$ (15)
where $\mu_{4}$ is the fourth central moment of the vorticity distribution $\mathcal{P}(\omega)$ and $\sigma$ is its standard deviation. The distribution of vorticities is sampled over times within a stationary configuration or a dynamical steady state, a few examples of which may be seen in Fig. 3(a). For a Gaussian distribution $K=0$, hence we use this as a measure of deviation from Gaussian statistics. In general, larger excess kurtosis indicates larger distribution tails or a sharper distribution peak. We note that excess kurtosis, and more generally vorticity distribution shape, has been previously used as a measure of active nematic phase behavior [55, 58, 61, 31].
To understand any ordering in the vorticity field, we also measure the correlation function of the vorticity field in either the $x$ and $y$ directions, given by
$C_{y}(x_{0},Y,t)$ $=\frac{\langle\omega(x_{0},y+Y,t)\omega(x_{0},y,t)\rangle_{y}}{\langle\omega(x_{0},y,t)^{2}\rangle_{y}}$ (17)
where $x_{0}$ or $y_{0}$ are fixed values and $\langle\cdot\rangle_{x,y}$ denotes an average over $x$ or $y$. Vortex ordering along these directions is identified with peaks in $C_{x,y}$, which we discuss further in Sec. IV.
III Turbulence Transitions
III.1 Activity Stripes
@@
Figure 2: (a-d) Spatial plots of (left) scaled time-averaged vorticity field $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ (color) with time-averaged velocity field $\langle\mathbf{v}\rangle$ (arrows) and (right) scaled time-averaged topological defect density $\langle D\rangle/|\langle D\rangle|_{max}$, where bright (dark) areas indicate $\alpha=\alpha_{0}$ ($\alpha=0$). The length of the black arrows indicate the relative magnitude of $\langle\mathbf{v}\rangle$. (a) $\ell_{i}=6$ and $\alpha_{0}=0.7$. (b) $\ell_{i}=16$ and $\alpha_{0}=0.7$. (c) $\ell_{i}=6$ and $\alpha_{0}=1.2$. (d) $\ell_{i}=16$ and $\alpha_{0}=1.5$. Animations of the dynamics of each state are available in the Supplemental Material [56].
We first focus on the qualitative transitions observed in the activity stripe system [Fig. 1(a)] upon varying $\ell_{i}$ and $\alpha_{0}$. For small $\ell_{i}$ and $\alpha_{0}$, the behavior resembles the active turbulence that would appear if the activity were homogeneous [4, 55, 6, 61]. We show an example of this in Fig. 2(a), where we plot the time-averaged normalized vorticity $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ and velocity $\langle\mathbf{v}\rangle$ fields as well as the time averaged defect density $\langle D\rangle/|\langle D\rangle|_{max}$ for a system with $\ell_{i}=6$ and $\alpha_{0}=0.7$. The average velocity and vorticity are close to zero throughout the domain, with no correlation between the active regions and the weak vortical patterns. Likewise, the topological defect density is also not correlated with the active regions, unlike in systems with a single activity stripe [34] or active channels [58].
When either $\alpha_{0}$ or $\ell_{i}$ is increased, as shown in Fig. 2(b) at $\ell_{i}=16$ and $\alpha_{0}=0.7$ and in Fig. 2(c) at $\ell_{i}=6$ and $\alpha_{0}=1.2$, $\langle\omega\rangle/|\langle\omega\rangle|_{max}$ becomes nonzero in the active regions while $|\langle\mathbf{v}\rangle|$ is suppressed in the active regions, which can be seen by the length of the black arrows in Fig. 2, so that steady flows occur only in the inactive regions between the active stripes. Additionally, the time-averaged topological defect density separates, and positive topological defects occupy the active regions while negative topological defects remain in the inactive regions. This charge separating behavior is similar to the behavior found in previous observations of active nematics in a single active stripe or confined in a channel [58, 34].
At large $\alpha_{0}$, the system develops stable and sometimes static vortices, as illustrated in Fig. 2(d) at $\ell_{i}=16$ and $\alpha_{0}=1.5$, where $\langle\mathbf{v}\rangle$ has clear vortical signatures. By plotting the time averages $\langle\omega\rangle$ and $\langle\mathbf{v}\rangle$ in this state versus those in Figs. 2(a,b,c), we see a clear difference in the long-time dynamical behavior of the system. The nematic configuration in this regime consists of two positive defects rotating around one another near the center of a given vortex, while negative defects sit in between the vortices. We discuss this state in more detail in Sec. IV.
Figure 3: (a) Vorticity distributions, $\mathcal{P}(\omega)$, for simulations with $(\ell_{i},\alpha_{0})=(6,0.7),\,(16,0.7),\,(6,1.2),\,(16,1.5)$ (the same parameter sets in Fig. 2). The dashed black line shows a Gaussian distribution. (b) Heat map of the logarithm of the excess kurtosis of the vorticity distribution over the whole domain, $\log|K|$, as a function of $\ell_{i}$ and $\alpha_{0}$. (c) Heat map of the excess kurtosis of the vorticity distribution measured over active regions only, $\log|K^{*}|$, as a function of $\ell_{i}$ and $\alpha_{0}$. Dashed line: points from panel (b) where $\log|K|=0$. Solid line: points from panel (c) where $\log|K^{*}|=0$. Letters correspond to parameter sets shown in Fig. 2.
To quantify the apparent transitions to and from active turbulence, we examine the vorticity distribution $\mathcal{P}(\omega)$ across the transitions. In Fig. 3(a) we plot examples of $\mathcal{P}(\omega)$ for the same parameter sets shown in Fig. 2. Only the simulation that qualitatively resembles two-dimensional active turbulence — $\ell_{i}=6$ and $\alpha_{0}=0.7$ — is characterized by a Gaussian vorticity distribution, shown by the black dashed line in Fig. 2(a). The other distributions are sharply peaked around $\omega=0$ and/or have long tails. We further characterize the distribution by the excess kurtosis of the distribution $K$, defined in Eq. (15). In Fig. 3(b) we plot a heat map of $\log|K|$ measured over the entire spatial domain as a function of $\alpha_{0}$ and $\ell_{i}$. The dashed line indicates points where $\log|K|=0$, and we define the $\log|K|<0$ region below this line to be in 2D active turbulence. We demarcate $K=1$ as the line for the 2D active turbulence transition because below this line the values of $K$ for each simulation do not vary much and tend to be close to $K=0$ which indicates a perfect Gaussian distribution. Additionally, we find that for systems characterized by vorticity distributions with $K>1$, the excess kurtosis increases with increasing $\alpha_{0}$ or $\ell_{i}$, while for systems with $K<1$ the small fluctuations in $K$ are not correlated with $\alpha_{0}$ or $\ell_{i}$.
@@
Above this transition, we observe that the qualitative behavior is similar to active systems confined to a channel or single activity stripe [55, 58, 34]. To probe this, we perform a measurement of $K[\mathcal{P}(\omega|\alpha>0.5\alpha_{0})]\equiv K^{*}$, in which we restrict the domain to only the active portions of the sample and exclude the inactive areas. We plot the resulting heat map of $\log|K^{*}|$ as a function of $\alpha_{0}$ versus $\ell_{i}$ in Fig. 2(c). The solid line indicates points for which $\log|K^{*}|=0$. In the window between the solid and dashed lines, where the active regions produce a turbulent signature but the inactive regions do not, 1D active turbulence has emerged along each active stripe, but turbulent behavior is suppressed in the regions between the stripes. Thus, the two transitions in the system are between 2D and 1D active turbulence, and between 1D turbulence and a vortex state.
III.2 Activity Circles